OpenStudy (anonymous):

for the data in the table does y vary directly with x if it does write an equation for the direct variation. x | y 10 12 15 18 20 24

OpenStudy (anonymous):

I will fan and medal!!

OpenStudy (janu16):

@mathwizzard3

OpenStudy (retireed):

You need to plot the data on a graph to try and visualize any correlation.

OpenStudy (mathwizzard3):

Watch this to get an understanding! https://www.youtube.com/watch?v=RLpmnRPtp0s

OpenStudy (dayakar):

X/y= k ( k is a constant ) K = x/y =10 /12 = 2×5/2×6=5/6 6x =5y

OpenStudy (mathwizzard3):

Please watch the video! It will help you A LOT!!!!!!!!!!

OpenStudy (mathwizzard3):

It is only 1 minute

OpenStudy (anonymous):

I watched the video. I kind of understand it now

OpenStudy (mathwizzard3):

Okay, so what do you think it is?

OpenStudy (mathwizzard3):

direct or indirect?

OpenStudy (dayakar):

There is no need of draw a graph to find a relation between x and y

OpenStudy (anonymous):

its direct = 1.2

OpenStudy (mathwizzard3):

Yes!

OpenStudy (anonymous):

thanks for sharing the video! i think I'm good with those kind of problems now.

OpenStudy (mathwizzard3):

That's great, I'm glad it helped!

OpenStudy (dayakar):

X = 5/6 y

OpenStudy (dayakar):

There is no need of draw a graph to find a relation between x and y

OpenStudy (dayakar):

X/y= k ( k is a constant ) K = x/y =10 /12 = 2×5/2×6=5/6 6x =5y

OpenStudy (mathwizzard3):

so is that all you need?

OpenStudy (anonymous):

yeah I might have another question later tho, can i tag you if I need help?

OpenStudy (mathwizzard3):

Okay but I might be leaving soon :)

OpenStudy (anonymous):

okay, I'm still working on the assignment so if I have a question it will be probably be soon anyway